A trailer truck traveling east at 24.0 m/s sounds a 600 Hz horn. What is the wavelength of the sound heard by a stationary listener located west of the truck? (The speed of sound is 336 m/s)
Answer #1

Given:

speed v = 24 m/s

frequency f = 600 Hz

speed of sound c = 336 m/s

Use the Doppler effect formula:

f’ = f * c/(c+v)

= 600 * 336/(336+24)

= 560 Hz

Thus the wavelength heard:

w = c/f’

= 336/560

= 0.6 m (answer)

Kindly upvote :)